OpenStudy (loser66):

How to use MVT to show for any n>0\(\dfrac{n}{2n+1}<\sqrt{n^2+n}-n <\dfrac{1}{2}\) Please, help

OpenStudy (loser66):

n >0, 2n+1>0, hence the far left >0, so that we have interval (0,1/2)

OpenStudy (loser66):

To apply MVT, we just show that \(\sqrt{n^2+n}-n >\dfrac{n}{2n+1}\) in (0,1/2)

OpenStudy (xapproachesinfinity):

i'm very short in time let's see

OpenStudy (xapproachesinfinity):

so did you show that part?

OpenStudy (xapproachesinfinity):

try to define a function where x>0 \[f(x)=\sqrt{x^2+x}-x-1/2\]

OpenStudy (xapproachesinfinity):

you need to show \[f(x)=f(0)+f'(c)x\leq0\]

OpenStudy (loser66):

I would like to know the middle part

OpenStudy (xapproachesinfinity):

that's for max min do another one

OpenStudy (loser66):

oh, got you.

OpenStudy (xapproachesinfinity):

that would show min<blah<max which is exactly what the MVT is about

OpenStudy (xapproachesinfinity):

there may be couple of other ways to do this but go for the simplest hehe

OpenStudy (xapproachesinfinity):

in your calc level they would only say few stuff and bam you got the result but that's advanced for me hehe

OpenStudy (loser66):

ok, thanks for help. Let me try.

OpenStudy (xapproachesinfinity):

ok! no problem i'm still here for a few mins let's see if that works

OpenStudy (loser66):

You can go, I need time to observe it and to arrange the stuff neatly. If I miss a small mistake, my credit is out. That is my prof's policy. hehehe

OpenStudy (loser66):

*make, not miss

OpenStudy (xapproachesinfinity):

haha that's a tough prof :) i would hate him for that, unfair haha

OpenStudy (loser66):

Nope, I love him, he trained me to be a good student who knows how to put the stuff flawless.

OpenStudy (xapproachesinfinity):

eh okay! may be he works for you after all :)

OpenStudy (loser66):

I got a C last semester with him. hehehe... but still love him.

OpenStudy (xapproachesinfinity):

haha i can't judge until i'm part of his lectures you know better than me :)

OpenStudy (xapproachesinfinity):

I like the kind of profs like MIT people you see the prof for calc he is awesome that is if you watch some mit videos

OpenStudy (xapproachesinfinity):

i let you finish this friend :) good luck

OpenStudy (loser66):

ty

OpenStudy (loser66):

I don't think it works on this way :( My attempt: Let \(f(x) = \sqrt{x^2+x}-x -\dfrac{1}{2}\) we need show \(f(x) = f(0) +f'(c) x < 0\) \(f(0) = -\dfrac{1}{2}\) and \(f'(x) =\dfrac{2x+1}{2\sqrt{x^2+x}}-1\) hence \(f(x) =-\dfrac{3}{2}+\dfrac{2c+1}{2\sqrt{c^2+c}}x\) This is an equation of a line with the slope >0 , it can't be <0 if x >0 

OpenStudy (xapproachesinfinity):

You made an error I guess

OpenStudy (loser66):

point out, please.

OpenStudy (xapproachesinfinity):

You didn't use 0<c<x for the interval Also how did you get - 3/2 check that again

OpenStudy (loser66):

Let me close and post it again. I messed it up. :)

OpenStudy (fibonaccichick666):

yea, you don't really want my help here loser, my googleing would be the only saving grace